question 3&4

question 3&4
media%2Fa0a%2Fa0a08e75-e09b-4339-a0ce-34
Answer #1

3.

orbital velocity is given by

Vo^2 = g(R + h)

Vo^2 = 1.04* ( 1700 + 472)*10^3

Vo = 1502.96 m/s

========

b)

T = 2 pi ( R + h) / Vo

T = 2* 3.14* (1700 + 472)*10^3 / 1502.96

T = 9072.53 s

========

4.

a = r w^2

2.9 = 70 *w^2

w = 0.2035 rad/s

w = 1. 94 rev/min

=======

Comment in case any doubt, will reply for sure.. Goodluck